A tube contains a 2m layer of olive oil floating above a 1.5m layer of water in turn floats above a 0.5m layer of mercury the upper layer is exposed to the pressure of the atmosphere of 1.01×10^5 N/m^2. What is the absolute pressure of the bottom of the mercury
1
Expert's answer
2018-06-07T07:45:08-0400
Pressure in this situation is additive, so pressure at the bottom of mercury layer will be a sum of atmosphere pressure, oil pressure and water pressure. p=p_atm+p_oil+p_water+p_mercury. Pressure of fluids can be defined as p=ρgh, where ρ – is a density of fluid, g – is standard acceleration of free fall and h – is thickness of a layer. So p=p_atm+ρ_oil gh_oil+ρ_water gh_water+ρ_mercury gh_mercury, where ρ_oil=900 kg/m^3 , ρ_water=1000 kg/m^3 , ρ_mercury=13500 kg/m^3 .
p=1.01×〖10〗^5+900×10×2+1000×10×1.5+13500×10×0.5==101000+18000+15000+67500=201500 N/m^2 ≈2×〖10〗^5 N/m^2